OpenStudy (anonymous):

f(x)=1+7/x-4/x^2 where is the function increasing/decreasing?

OpenStudy (anonymous):

find first derivative. Later look for intervals of x axis where this derivative is negative/positive, since that means that function is decreasing/increasing there

OpenStudy (anonymous):

for the derivative i get \[\frac{ -7x+8 }{ x^3 }\] so x=8/7

OpenStudy (anonymous):

there is a vertical asymptote at 1 and a horizontal at 0

OpenStudy (anonymous):

\[f'(x)=-\frac{7}{x^{2}}+\frac{8}{x^{3}}\] this is not defined at x=0. So you hae to study what happens for x>0 and x<0 1º Let's start with x>0. f'(x)=0 at 7/x^2=8/x^3 => x=8/7 for x>0 and x<8/7 f'(x)>0. so function increase for x>0 and x>8/7 f'(x)<0. So function decrease 2º x<0 In this case f'(x)<0 for all x. So function decrease

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

so it would be increasing:(0,8/7) decreasing: (8/7, inf)?

OpenStudy (anonymous):

no: increasing:(0,8/7) decreasing: (-inf,0) union (8/7, inf)

OpenStudy (anonymous):

@guest1234